OpenStudy (anonymous):

Write each fraction in terms of the LCM of the denominators. x+1 over x+5 x+3 over x-8 4 over 40+3x-x2

OpenStudy (anonymous):

what doest it mean in "terms of the LCM of the denominators."?

OpenStudy (anonymous):

40+3x-x^2 = -(x-8) (5+x) x-8 x+5 40+3x-x^2 is the LCM

jhonyy9 (jhonyy9):

so i think that this sign low common denominator and in case of those fravction i think that will be x2-3x-40=(x-8)(x+5)